OpenStudy (anonymous):

Cylinder A has a radius of 12 inches and a height of 6 inches. Cylinder B has a volume of 648π. What is the percent change in volume between cylinders A and B? Cylinder B is 50% smaller than cylinder A. Cylinder B is 25% smaller than cylinder A. Cylinder B is 150% bigger than cylinder A. Cylinder B is 200% bigger than cylinder A.

OpenStudy (anonymous):

@Hero @.Sam. can you help me?

OpenStudy (anonymous):

@lalaly

hero (hero):

Hint: Use the volume of a cylinder formula and the percent change formula to figure it out.

hero (hero):

Volume of a Cylinder: \[V_c = \pi r^2 h\] Percent change: \[\frac{\%\Delta}{100} = \frac{|a - b|}{a}\]

OpenStudy (anonymous):

π(12)^2(6) right?

OpenStudy (anonymous):

i got 275.16

hero (hero):

Looks right. Finish the calculation

OpenStudy (anonymous):

how would i do that?

hero (hero):

Use the "percent change" formula

OpenStudy (anonymous):

im sorry. i am so confused